OpenStudy (anonymous):

What is the perimeter of a rectangle with a length of 17 and a width of 11?

OpenStudy (anonymous):

\[17+17+11+11\] \[2\times 17+2\times 11\] \[2(17+11)\] your choice

OpenStudy (anonymous):

Perimeter = 2 x length + 2 x width. Perimeter = 2(17) + 2(11) = 34 + 22 = 56